OpenStudy (anonymous):

(3,4) is the midpont of JK , where J . has coordinates (1,2)

OpenStudy (anonymous):

So you need K ?

OpenStudy (carson889):

X: 3 = (1+ x)/2 Y: 4 = (2+y)/2 X: 6 = 1+x Y: 8 = 2+y X: 6-1 = x = 5 Y: 8-2 = y = 6 K = (5,6)

OpenStudy (anonymous):

I got (2,3) what did I do wrong?

OpenStudy (carson889):

Since the formula for midpoint is: \[\left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 } \right)\] = (midpoint x, midpoint y) We know x1 and y1 are the values of J, and the midpoints are x = 3, y = 4. x2 and y2 are the values of K \[\left( 3=\frac{ 1+x _{2} }{2 },4 = \frac{ 2+y _{2} }{ 2 } \right)\] \[ \left( 6=1+x_{2},8=2+y _{2} \right)\]